OpenStudy (anonymous):

Solve 3x^3 - 4x^2 - 12x + 16 = 0

OpenStudy (anonymous):

we can write: (3x^3 - 12x) + (-4x^2 +16) = 0 3x^2(x-4)-4x(x-4)=0

OpenStudy (anonymous):

What are you wanting to do exactly?

OpenStudy (anonymous):

There are several ways to answer this equation

hero (hero):

(3x^3 - 12x) + (-4x^2 +16) = 0 3x^2(x-4)-4x(x-4)=0 (x-4)(3x^2-4x) = 0 x(3x-4)(x-4) = 0 x = 0, 4/3, 4

OpenStudy (anonymous):

thanks. I forgot minus symbol

hero (hero):

I copied what you already wrote and just finished it